# On a certain subclass of strongly starlike functions

### v3

@@ 3. Main Results
Let $\widetilde{G}$ be of the form (3.11). If $f(z)\in\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$, then
$r\exp\widetilde{G}(-r)<\left|f(z)\right|<r\exp\widetilde{G}(r),$
for each $r=|z|<1$.
@@
Let $0<\alpha_{1},\alpha_{2}\leq 1$, $c=e^{\pi i\theta}$ and $\theta=(\alpha_{2}-\alpha_{1})/(\alpha_{2}+\alpha_{1})$. Let $f\in\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$ and the coefficients of $\log(f(z)/z)$ be given by (3.9). Then
(3.15) $|\gamma_{n}|\leq\frac{(\alpha_{1}+\alpha_{2})}{2n}\cos\frac{\pi}{2}\theta.$
The result is sharp.
@@
$2n|\gamma_{n}|\leq|A_{1}|,$
where
$A_{1}=\frac{(1+c)(\alpha_{1}+\alpha_{2})}{2}.$
Thus, the desired inequality (3.15) follows. The equality holds for the logarithmic coefficients of the function
@@
If we let $\alpha_{1}=\alpha_{2}=\beta$ in the above Theorem 3.7, we get the following result, which previously was obtained by Elhosh, see [9, Theorem 2].
@@
$|\gamma_{n}|\leq\frac{1}{n}\beta.$
@@
We end this paper with the following result:
@@
${\rm Re}\left\{\frac{zf^{\prime}(z)}{f(z)}\right\}\geq\left(\frac{1-\left(1+2\cos(\pi\theta/2)\right)r}{1-r}\right)^{(\alpha_{1}+\alpha_{2})/2},\quad 0\leq|z|=r\leq\frac{1}{1+2\cos(\pi\theta/2)}$
@@
$0<{\rm Re}\left\{\frac{zf^{\prime}(z)}{f(z)}\right\}\leq\left(\frac{1+\left(2\cos(\pi\theta/2)-1\right)r}{1-r}\right)^{(\alpha_{1}+\alpha_{2})/2},\quad 0\leq|z|=r<1,$
where $0<\alpha_{1},\alpha_{2}\leq 1$ and $\theta=(\alpha_{2}-\alpha_{1})/(\alpha_{2}+\alpha_{1})$.
@@
$\left|F(z)-1\right|=\left|\frac{(1+c)w(z)}{1-w(z)}\right|\leq\frac{2|w(z)|\cos(\pi\theta/2)}{1-|w(z)|}\leq\frac{2r\cos(\pi\theta/2)}{1-r},\quad(|z|=r<1),$
which completes the proof of Theorem 3.8. ∎
@@
$\left(\frac{1-3r}{1-r}\right)^{\beta}\leq{\rm Re}\left\{\frac{zf^{\prime}(z)}{f(z)}\right\}\leq\left(\frac{1+r}{1-r}\right)^{\beta}\qquad(|z|=r\leq 1/3).$

### v4

@@ 3. Main Results
Let $\widetilde{G}$ be of the form (3.15) and $0<\alpha_{1},\alpha_{2}\leq 1$. If $f(z)\in\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$, then
$r\exp\widetilde{G}(-r)\leq\left|f(z)\right|\leq r\exp\widetilde{G}(r),$
for each $r=|z|<1$. Both inequalities are sharp.
@@
Let $f\in\mathcal{S}^{*}(\alpha_{1},\alpha_{2})$ and the coefficients of $\log(f(z)/z)$ be given by (3.13) and $0<\alpha_{1},\alpha_{2}\leq 1$. Then
(3.19) $|\gamma_{n}|\leq\frac{\lambda}{2n},\quad(n\geq 1),$
where $\lambda$ is given by (2.7). The result is sharp.
@@
$2n|\gamma_{n}|\leq|A_{1}|=\lambda.$
Thus, the desired inequality (3.19) follows. Equality holds for the logarithmic coefficients of the function
@@
If we let $\alpha_{1}=\alpha_{2}=\beta$ in the above Theorem 3.7, we get the following result, which was previously obtained by Elhosh, see [9, Theorem 2].
@@
$|\gamma_{n}|\leq\frac{1}{n}\beta,\quad(n\geq 1).$
@@
3.2. Upper and Lower Bounds for ${\rm Re}\{zf^{\prime}(z)/f(z)\}$, where $\in S^{*}(\alpha_{1},\alpha_{2})$
We finish this paper by estimating ${\rm Re}\{zf^{\prime}(z)/f(z)\}$, where $f\in S^{*}(\alpha_{1},\alpha_{2})$.
@@
$\mathrm{Re}\left\{\frac{zf^{\prime}(z)}{f(z)}\right\}\geq\left(\frac{1-\left(1+2\cos(\pi\theta/2)\right)r}{1-r}\right)^{(\alpha_{1}+\alpha_{2})/2},\quad 0\leq|z|=r\leq\frac{1}{1+2\cos(\pi\theta/2)},$
@@
$\mathrm{Re}\left\{\frac{zf^{\prime}(z)}{f(z)}\right\}\leq\left(\frac{1+\left(2\cos(\pi\theta/2)-1\right)r}{1-r}\right)^{(\alpha_{1}+\alpha_{2})/2},\quad 0\leq|z|=r<1,$
where $0<\alpha_{1},\alpha_{2}\leq 1$ and $\theta=(\alpha_{2}-\alpha_{1})/(\alpha_{2}+\alpha_{1})$. Moreover, $\mathrm{Re}\{zf^{\prime}(z)/f(z)\}>0$ for all $z\in\mathbb{D}\setminus\{0\}$.
@@
$\left|F(z)-1\right|=\left|\frac{(1+c)w(z)}{1-w(z)}\right|\leq\frac{2|w(z)|\cos(\pi\theta/2)}{1-|w(z)|}\leq\frac{2r\cos(\pi\theta/2)}{1-r},\quad(|z|=r<1).$
This inequality implies that
$\frac{1-(1+2\cos(\pi\theta/2))r}{1-r}\leq\mathrm{Re}\,\{F(r)\}\leq\frac{1+(2\cos(\pi\theta/2)-1)r}{1-r}.$
For functions with positive real part, raising to the power $(\alpha_{1}+\alpha_{2})/2$ preserves these inequalities, giving the stated bounds. Of course, the constraint
$r\leq\frac{1}{1+2\cos(\pi\theta/2)}$
ensures the lower bound remains non-negative. This concludes the proof of Theorem 3.8. ∎
@@
$\left(\frac{1-3r}{1-r}\right)^{\beta}\leq{\rm Re}\left\{\frac{zf^{\prime}(z)}{f(z)}\right\}\leq\left(\frac{1+r}{1-r}\right)^{\beta},\quad(|z|=r\leq 1/3).$